# An analytical model for the remote epitaxial potential

### v1

@@ Abstract
We propose an analytical model for the remote bonding potential of the substrate that permeates through graphene during remote epitaxy. Our model, based on a Morse interatomic potential, includes the attenuation due to (1) the increased separation between film and substrate and (2) free carrier screening from graphene. Compared with previous slab density functional theory calculations, which use the electrostatic potential as a proxy for bonding, our analytical model provides a more direct description of bonding, explicitly includes screening (which is often ignored), and is based on simple, interpretable, and well benchmarked parameters. We show that the magnitude of for typical semiconductor and oxide substrates is few meV, similar to the van der Waals potential of graphene. This suggests interference between the graphene and remote substrate potentials, plus interfacial reconstructions, must be considered when interpreting experiments on remote epitaxy. We use our model to interpret previous experiments from the remote epitaxy and related literature, highlighting connections to moire epitaxy and to the ordering of graphite intercalation compounds. Our model also points to tests, based on tunable screening and spatial extent of the substrate potential, that may increase the strength of the remote potential towards the more idealized picture.
@@
However, there are several challenges to this understanding. First, it is not clear that a purely electrostatic interaction can explain remote epitaxy, since the electrostatic potential describes the electronic workfunction rather than bonding (i.e., the energy of pulling an electron to infinity from that location). Electrostatic potentials are thus completely agnostic to the chemical match between the film and substrate, which has been shown to impact remote epitaxy efficacy [9]. Second, most DFT treatments ignore free carrier screening from graphene, ignore the lattice potential of graphene, and ignore possible reconstructions. Finally, DFT calculations can be a black box, making it difficult to interpret the contributions from various terms without additional tests. A simple analytical model is needed in order to build a physical understanding of the various terms that contribute to the strength of the remote potential.
@@
Here we develop an analytical model for the remote bonding potential permeation through graphene. Our model is based on an empirical Morse interatomic potential [10], which includes contributions from both covalent bonding and electrostatic interactions [11]. This simple two body potential captures the essential of physics and employs easy to understand parameters that are well benchmarked by experiments, density functional theory, and molecular dynamics simulations [12]. Compared to previous approaches that used the electrostatic potential, our approach (1) uses a potential that includes both covalent bonding and electrostatic interactions [11], (2) explicitly includes free carrier screening of graphene, which is often ignored [6, 1, 8], (3) avoids the use of an adjustable scale factor that has been invoked to make the electrostatic potential behave more like a bonding potential [2], and (4) avoids the use of complicated background subtraction and data cropping [2]. We show that graphene is not as transparent as often assumed: the remote potential of the substrate through graphene has an order of magnitude few meV or smaller, similar to the the vdW interaction strength of graphene itself. Our model predicts that the spatial extent of the bond potential, rather than bond ionicity, is the primary parameter that controls the strength of the remote potential.
Our model treats graphene in ideal limit as a homogeneous electron gas. The key elements are (1) increased separation between film and substrate due to insertion of the graphene and (2) attenuation due to free carrier screening. We start from a Morse potential for the substrate
@@
where the first term describes the short-range Pauli repulsion and the second term is the long range attraction. Here is the bond strength, is the width of the potential well, and is the equilibrium bond distance. The potential for a bare GaAs substrate without a remote spacer layer is shown in Fig. 1(a) (black curve). This potential describes the pair bonding interaction between an As atom and a Ga atom, and thus is appropriate for homoepitaxy of a GaAs film on a GaAs substrate. We use parameters eV and Å-1, based on optical spectroscopy measurements of the bond strength and oscillation frequency of GaAs [13], and refined by ab initio calculations in a many-body form to reproduce the bulk properties [14, 15]. We let Å, corresponding to a generalized layer spacing between atomic planes.
@@
where we let meV and Å. The resulting potential , in the absence of screening, is shown in the blue curve. The inert spacer displays a new minimum of meV, which is only of the Ga-As bond energy. The new equilibrium film-substrate separation of Å is within the 5-6 Å separation typically observed by TEM of GaAs/graphene/GaAs [16]. Our insulating spacer calculation is consistent with previous slab DFT calculations that use charge density as a proxy for the remote potential permeation: for (001)-oriented GaAs slabs separated by a 7 Å vacuum gap (no graphene for screening), the computed charge density in the middle of the gap was of the charge density in the interior of the material [1].
@@
where Å is the effective thickness of graphene and is the number of graphene layers. Figure 1(c) plots through monolayer and bilayer graphene, as a function of typical room temperature graphene carrier densities cm-2 (). We find that the screening is highly tunable with carrier density, and there is a significant screening-induced attenuation going from monolayer graphene to bilayer graphene (in addition to the attenuation from increased layer separation). For greater than the thickness of graphene, the total screened potential is
where the first term is the screened remote potential of the substrate and the second term is the Lennard-Jones (vdW) potential of graphene.
@@
At first glance this may seem inconsistent with previous DFT slab calculations of the permeated potential, which suggested 15 meV fluctuations through graphene on GaAs substrates [2]. However, we emphasize that (1) the previous calculation was for the electrostatic potential [2], which is not the bonding potential, (2) the calculations employed selective cropping and replication and thus is not a reliable representation of the potential fluctuations [2], and (3) in many calculations the screening effects from graphene and the lattice potential of graphene itself are not explicitly included.
@@
We now apply our model to other commonly used semiconductor and oxide substrates. Table 1 summarizes the Morse potential parameters for ZnO [18], GaN [19], GaAs [14, 15], SiC [20], Si [20], and Cu [21]. We first use these parameters to estimate the attenuation from the separation induced by the offset of an inert 2D materials spacer, compared to the attenuation from screening. For simplicity, we use the fact that for large the attractive term of the Morse potential (Eqn 1) dominates. We thus use . In Table 1 we compute at Å. We find that the separation induced attenuation for these substrates of to is slightly more significant than the screening induced for monolayer graphene at typical carrier densities cm-2 (Fig. 1(b)).
We next test the hypothesis from Ref. [2] that polar ionic bonding controls the strength of the remote potential. Table 1 summarizes the remote potential under the approximation
@@
Interestingly, our model suggests that monoatomic Si should have the strongest remote potential among the substrates considered. This lies in apparent contradiction to the experiments in Ref. [2] reported that growth of Si on graphene/Si (001) resulting in a polycrystalline film, yet growth on substrates with more polar bonding enabled epitaxial films in more ionic LiF substrates that were covered with up to three layers of graphene. However, the impact of key synthesis parameters including substrate temperature, fluxes, type of precursor (e.g., atom, molecular cluster, metalorganic), and graphene surface preparation remain unclear.
@@
We caution that our model should only be considered an order of magnitude estimate for the remote potential. We treat graphene as an idealized homogeneous electron gas, and only consider pair-wise atom-atom interactions between film and substrate. Our model does not consider the directional character of bonding, multi-atom interactions, or bonding between the film and graphene or graphene and the substrate. It also does not consider surface reconstructions. Still, it identifies an essential role for the decay with distance and for screening. These essential parameters place significant constraints on the magnitude of the remote bonding potential that are overlooked in the current atomistic calculations. Experimental techniques such as He atom scattering [22] and force-distance spectra by atomic force microscopy [23, 24, 25, 26], as a function of graphene gating, could directly measure the remote potential as a function of screening to test our predictions.
Given the relatively weak remote potential, it is worth considering other explanations for remote epitaxy, beyond the simple picture of transparent graphene. For the bulk graphite intercalations compounds, the intercalants form ordered superlattices that can be commensurate or incommensurate to the graphite lattice [27]. This intercalant ordering is thought to be governed by several factors, including (1) Coulomb repulsion between intercalants, since there is charge transfer between the intercalants (often alkali metals) and the graphite sheets, and (2) strain fields in the graphite sheets induced by bending around the intercalants [28]. These factors may also be relevant to remote epitaxy. The former charge transfer is not a remote interaction per se, but hints at a possible extended molecular orbital picture that includes atomic orbitals from the film and substrate, hybridized with the extended system of graphene. Such a description is not included in our model. The second factor, strain and buckling in graphite, suggests that surface reconstructions may be relevant to epitaxy on graphene-covered substrates. Indeed, strong buckling is observed in the “buffer” layer of graphene that forms on Si-face SiC (0001) [29]. This distorted graphite has a mix of and bonds. The out of plane dangling bonds are expected produce bonding potential fluctuations of order 100 meV to eV (since they are covalent bonds), larger than the few meV scale remote and vdW potentials. Indeed, recent experiments demonstrate the adatom sticking coefficient is higher on buffer graphene than on epitaxial graphene [30], implying higher reactivity due to these dangling bonds. These reconstructions and moire interference with the graphene are known to template the ordering of metal adatoms on graphene/Ir (111) [31], graphene/Rh (111) [32], graphene/Pt (111) [33], graphene/SiC (0001) [34]. Atomistic calculations that include these reconstructions as well as direct chemical interaction between the film and substrate are required to fully understand remote epitaxy of continuous films on graphene.

### v2

# Analytical model for the remote epitaxial potential

@@ Abstract
We propose an analytical model for the remote bonding potential of the substrate that permeates through graphene during remote epitaxy. Our model, based on a Morse interatomic potential, includes the attenuation due to the increased film-substrate separation and due to graphene free carrier screening. Compared with previous slab density functional theory calculations, which use the electrostatic potential as a proxy for bonding, our analytical model includes covalent and van der Waals bonding interactions, includes screening (which is often ignored), and is based on simple, physically interpretable, and well benchmarked parameters that build understanding. We show that for typical graphene free carrier concentrations of order cm-2, the magnitude of for most semiconductor and oxide substrates is few to tens of meV, similar to the van der Waals potential of graphene. This suggests interference between the graphene and remote substrate potentials must be considered when interpreting experiments on remote epitaxy. Furthermore, we show that (1) the strength of the remote potential is tunable by screening (graphene carrier density), (2) the remote potential is vanishingly weak through two or more graphene layers, (3) the spatial extent of the potential, rather than degree of ionicity, controls the strength of the permeated bonding potential. Our model provides a simple framework for benchmarking direct measurements of the remote epitaxial potential, and we propose several experimental paths to measure this quantity. De-convolving the effects of the native remote potential, graphene defects, and tunable growth kinetics is key to understanding and tuning the mechanisms of remote epitaxy.
@@
However, there are several challenges to this understanding. First, DFT slab calculations can be a computationally expensive black box, making it difficult to interpret the contributions from various terms without additional tests. A simple analytical model is needed to build a physical understanding of the various terms that contribute to the strength of the remote potential, and enable rapid surveys over candidate materials. Second, although electrostatics define important electronic quantities like the workfunction and band alignments, the formation and structure of crystals are instead typically driven by covalent and van der Waals (vdW) bonding interactions between the substrate and film. New models need to consider these covalent and vdW interactions, rather than the pure electrostatics focus of existing DFT, which often does not incorporate any knowledge of the film.
Finally, theories of the remote potential are difficult to benchmark because there are few direct measurements of the remote potential [9, 10]. Instead, its strength is typically inferred from the final outcomes of growth experiments [2, 1, 6, 8], which convolve the remote potential with defects and growth kinetics [11]. Defects and kinetics have a significant impact on growth mechanisms: previous experiments [12] demonstrate that since typical adatom diffusion lengths on graphene are 100 nm to microns at typical growth temperatures [13, 14, 12, 15], pinhole-seeded epitaxy, rather than remote epitaxy, dominates if the defect density on graphene exceeds 1 [12, 11]. Since most experiments do not report the graphene defect density after the critical anneal step that produces pinholes [12], most experiments to date cannot distinguish remote epitaxy from pinhole epitaxy. An analytical theory of the remote potential, with physically interpretable parameters that can be benchmarked against direct measurements of the remote potential, is required to de-convolve the various factors that dictate the growth mechanisms on graphene-covered surfaces.
@@
Here we develop an analytical model for the remote bonding potential permeation through graphene. Compared to previous DFT-based electrostatic potential calculations, our model (1) is based on simple physical parameters to aid physical understanding rather than black box approaches, (2) includes more physically motivated covalent bonding and vdW interactions between the substrate and film rather than purely electrotatics [16], and (3) includes free carrier screening of graphene, which is often ignored [6, 1, 8]. We show that graphene is not as transparent as often assumed: for typical graphene carrier densities, the remote potential of the substrate through graphene has an order of magnitude few meV, similar to the the vdW interaction strength of graphene itself. Our model predicts that the spatial extent of the bond potential, rather than bond ionicity, is the primary parameter that controls the strength of the remote potential. Finally, our model identifies graphene screening as an attractive parameter for continuously tuning the strength of the remote potential. We propose direct measurements of the remote potential based on scanning probe and atom scattering techniques to benchmark our model, as a function of gate-tuned graphene carrier density.
We start from an empirical Morse interatomic potential [17], which includes covalent, vdW, and electrostatic interactions [16], to describe the substrate. This simple two body potential captures the essential physics and employs easy to understand parameters that are well benchmarked by experiments, density functional theory, and molecular dynamics simulations [18], before the inclusion of graphene. The Morse potential has the form
@@
where the first term describes the short-range Pauli repulsion and the second term is the long range attraction. Here is the bond strength, is the width of the potential well, and is the equilibrium bond distance. The potential for a bare GaAs substrate without a remote spacer layer is shown in Fig. 1(a) (black curve). This potential describes the pair bonding interaction between an As atom and a Ga atom, and thus is most appropriate for homoepitaxy of a GaAs film on a GaAs substrate (or to a limited extent, heteroepitaxy with Ga-As interface termination, e.g. InAs/GaSb (001)). We use parameters eV and Å-1, based on optical spectroscopy measurements of the bond strength and oscillation frequency of GaAs [19], and refined by ab initio calculations in a many-body form to reproduce the bulk properties [20, 21]. We let Å, corresponding to a generalized layer spacing between atomic planes.
@@
where we let meV, Å, and is the equilibrium position of the new 2D material spacer. The total potential , in the absence of screening, is shown in the blue curve. The inert spacer displays a new minimum of meV, which is only of the Ga-As bond energy. The new equilibrium film-substrate separation of Å is within the 5-6 Å separation typically observed by TEM of GaAs/graphene/GaAs [22]. Our insulating spacer calculation is consistent with previous slab DFT calculations that use charge density as a proxy for the remote potential permeation: for (001)-oriented GaAs slabs separated by a 7 Å vacuum gap (no graphene for screening), the computed charge density in the middle of the gap was of the charge density in the interior of the material [1].
@@
where Å is the effective thickness of graphene and is the number of graphene layers. Figure 1(c) plots through monolayer and bilayer graphene, as a function of typical room temperature graphene carrier densities cm-2 (). We find that the screening is highly tunable with carrier density, and there is a significant screening-induced attenuation going from monolayer graphene to bilayer graphene (in addition to the attenuation from increased layer separation).
For greater than the thickness of graphene, the total screened potential is
where the first term is the screened remote potential of the substrate and the second term is the Lennard-Jones (vdW) potential of graphene. Remote epitaxy is expected if , i.e. if the screened substrate potential dominates over the graphene potential. This would produce an epitaxial film aligned to the substrate. Van der Waals epitaxy is expected if , resulting in a film that grow in the lowest surface energy orientation, with a weak driving force for in-plane alignment.
@@
At first glance this may seem inconsistent with previous DFT slab calculations of the permeated potential, which suggested 15 meV fluctuations through graphene on GaAs substrates [2]. However, we emphasize that (1) the previous calculation was for the electrostatic potential [2], which is not the bonding potential, (2) the calculations in Ref. [2] employed selective cropping and replication and thus are not a reliable representation of the potential fluctuations [11], and (3) in many calculations the screening effects from graphene and the lattice potential of graphene itself are not explicitly included.
@@
We now apply our model to other commonly used semiconductor and oxide substrates. Table 1 summarizes the Morse potential parameters for ZnO [23], GaN [24], GaAs [20, 21], SiC [25], Si [25], Ge [26], and Cu [27]. We first use these parameters to estimate the attenuation from the separation induced by the offset of an inert 2D materials spacer, compared to the attenuation from screening. For simplicity, we use the fact that for large the attractive term of the Morse potential (Eqn 1) dominates. We thus use . In Table 1 we compute at Å. We find that the separation induced attenuation for these substrates of to is slightly more significant than the screening induced for monolayer graphene at typical carrier densities cm-2 (Fig. 1(b)).
We next test the hypothesis from Ref. [2] that polar ionic bonding controls the strength of the remote potential. For a fixed graphene carrier density of cm-2, Table 1 summarizes the remote potential under the approximation
@@
Interestingly, our model suggests that monoatomic Si should have the strongest remote potential among the substrates considered. This lies in apparent contradiction to the experiments in Ref. [2] reported that growth of Si on graphene/Si (001) resulting in a polycrystalline film, yet growth on substrates with more polar bonding enabled epitaxial films in more ionic LiF substrates that were covered with up to three layers of graphene. However, the impact of key synthesis parameters including substrate temperature, fluxes, type of precursor (e.g., atom, molecular cluster, metalorganic), and graphene surface preparation remain unclear. We also note that in this case, carbon in the graphene may bond with the silicon substrate, altering the graphene potential from the straightforward screening assumption of our model.
@@
We caution that our model should only be considered an order of magnitude estimate for the remote potential. We treat graphene as an idealized homogeneous electron gas, and only consider pair-wise atom-atom interactions between film and substrate. Our model does not consider the directional character of bonding, multi-atom interactions, or bonding between the film and graphene or graphene and the substrate. It also does not consider surface reconstructions. For lattice mismatched heteroepitaxy, it does not consider strain or misfit dislocation energies. Still, it identifies an essential role for the decay with distance and for screening. These essential parameters place significant constraints on the magnitude of the remote bonding potential that are overlooked in the current atomistic calculations. More accurate state-of-the-art bond potentials, e.g. Tersoff potentials used in molecular dynamics simulations, could be applied to our framework of Eq. 4 to enhance our the model’s predictive power. Note, however, that more complex yet accurate potentials come at the cost of losing simple physically interpretable parameters. Additionally, future work could move Eq. 4 from a 1D model of bonding to a 2D model which could allow for the incorporation of additional physics such as reconstructions.
Our calculations also highlight the need for direct measurements of the remote potential. Current experiments rely on the final outcomes of growth to infer the strength of the remote potential through graphene. However, such experiments convolve the impact of the remote potential with defects and variable growth kinetics, thus obscuring an understanding of the growth mechanisms. Our model, which contains physically understandable parameters of potential width , potential depth , and screening in an analytical form, facilitates comparisons with direct measurements to benchmark our theory. We propose frequency-shift atomic force microscopy [28, 29, 30, 31] and He atom scattering [32] as methods to directly measure the covalent and vdW contributions, to complement the Kelvin-probe force microscopy [10] and photoemission/low energy electron microscopy (PEEM/LEEM) [9] that have mapped the electrostatic potential (workfunction) above graphene-covered substrates. These measurements could be performed as a function of gate-induced graphene screening to test our predictions of Eq 4. DFT calculations could also be used to computationally examine these remote potentials, provided they fully include any substrate-graphene reconstruction, the spacer layer of graphene, and at least one atom of the film layer to measure full chemical bonding and not just electrostatic potential.